5 pounds of caramel bats, costing 2.20 per pound are used.10 pounds of licorice bromstick, costing 1.80 per pound are also used. How many pounds of marshmallow goblins ($3 per pound) are needed if the candy mix sells for 2.50 per pound.
Answer by mananth(16946)

5 pounds of caramel bats, costing 2.20 per pound
10 pounds of licorice bromstick, 1.80 per pound
x pounds marshmallow goblins ($3 per pound)
candy mix sells for 2.50 per pound.15+x pounds
...
5*2.20+10*1.8+3x= 2.50 ( 15+x)
11+18+3x =37.5 +2.50x
3x-2.50x = 37.5-29
x= 8.5 pounds marshmallow goblins
...
m.ananth@hotmail.ca